User Tools

Site Tools


operations:agenda2016_04_21

Differences

This shows you the differences between two versions of the page.

Link to this comparison view

Both sides previous revision Previous revision
Next revision
Previous revision
Last revision Both sides next revision
operations:agenda2016_04_21 [2016/04/20 12:40]
Ellen Manning
operations:agenda2016_04_21 [2016/04/21 02:16]
Ellen Manning
Line 9: Line 9:
 Current Issues Current Issues
  
-    * New PCFS at Hobart (thanks Jamie)+    ​* Problems with drive hydraulics on 26m, cancelled today'​s experiment (RD1603). 
 +       - New something being bought by Eric 
 + 
 +    ​* New PCFS at Hobart ​is working great (thanks Jamie). Has anyone had any problems?
  
     * Maser maintenance at both Katherine and Yarragadee currently, hence Brett is not at the Observatory. Warren is at Mt Pleasant Tuesdays to Thursdays and occasionally Mondays.     * Maser maintenance at both Katherine and Yarragadee currently, hence Brett is not at the Observatory. Warren is at Mt Pleasant Tuesdays to Thursdays and occasionally Mondays.
  
-    * CRDS82 - "still getting ve error: ERROR ch -308 ve total power integrator overflow"​ - what was this and was it fixed Bryn?+    * CRDS82 - "still getting ve error: ERROR ch -308 ve total power integrator overflow"​ - Bryn, what was this and was it fixed? 
 +      - High tsys values due to RFI
  
     * R4734 - Yarragadee: consistent offset between mark 5 and other computers of 0.1 - 0.4 visible in fmset with no sync error, could not be fixed. Caused by power failure affecting GPS receiver (new UPS installed) - had to change to Hobart'​s. Fixed itself 4+ hours later. Is there any way to fix this other than hope?     * R4734 - Yarragadee: consistent offset between mark 5 and other computers of 0.1 - 0.4 visible in fmset with no sync error, could not be fixed. Caused by power failure affecting GPS receiver (new UPS installed) - had to change to Hobart'​s. Fixed itself 4+ hours later. Is there any way to fix this other than hope?
 +     - Probably ntp problem
  
     * R1735 - A couple of instances where recorder started ~ 7-10 seconds before onsource at both Katherine and Yarragadee.     * R1735 - A couple of instances where recorder started ~ 7-10 seconds before onsource at both Katherine and Yarragadee.
 +       - possibly due to tag along experiment
 +
 +    * High system temperatures - run systemp12 again to make sure it wasn't just one dodgy reading. Monitor and make note of how many scans were affected by consistent high tsys values.
 +      - elevation and weather affects tsys values
 +      - would be cool if monica can track changes
  
-    * High system temperatures - run systemp12 again to make sure it wasn't just one dodgy reading. Monitor make note of how many scans were affected by high tsys values.+    * Anything else to note
 +    * VSNs all reported as the same for 4 telescopes in end message (Jesse check experiment) - keep an eye on this
  
/home/www/auscope/opswiki/data/pages/operations/agenda2016_04_21.txt · Last modified: 2016/04/21 02:17 by Ellen Manning